Suppose assets are not perfectly substitutable. In particular, suppose that the domestic country must pay a risk premium given by p=0.0004(B-A), where B denotes the stock of domestic government debt held by the public (not central bank) and A denotes domestic assets held by central bank. Assume that the domestic real demand for liquidity is L(R,Y)= 2Y/500R, the domestic income is 1000 and foreign interest rate is 10%. Assume also that B=900, A=500 and Ms =1000. Finally, suppose that the nominal exchange rate is E=2.0 and that P=1.5

Question
1) What is the domestic interest rate R?(answer with a number instead of a percentage)
2)What is the risk premium for the domestic currency?(Answer with a number instead of a percentage)
3)If E=2, what is the equilibrium exchange rate?
Answer #1

Part 1

At equilibrium,

            Money Supply = Money demand

As per question,

            Money Supply = 1000

            Money demand = 2Y/500R

            Domestic income Y = 1000

Putting them in the 1st equation and solving for domestic interest rate R,

            1000 = (2 * 1000) / (500 * R)

            R = (2*1000)/(500 * 1000) = 0.004

Part 2

Given that,

B=900

A=500

As per the question, the domestic country must pay a risk premium given by,

p=0.0004(B-A) = 0.0004(900-500) = 0.0004*400 = 0.16

Part 3

Given that,

            Domestic interest rate rdom =.004

            Foreign interest rate rfor = .1

            Nominal exchange rate e = 2

Let E* be equilibrium exchange rate. By interest parity,

e* = e ( rfor / rdom) = 2 * ( 0.1 / 0.004 ) = 50